# Dark tranquility: Overdamped quasibound states inside a Schwarzschild black hole

### v1

@@ Abstract
Schwarzschild black-hole interiors, bounded by event horizons and terminated by spacelike singularities, are regions where all physical observers are inevitably destroyed. However, the behavior of wave propagation in such extreme habitats remains rich and nuanced. In this work, we show that axial gravitational perturbations in the interior of a Schwarzschild black hole can form overdamped (non-oscillatory) quasibound states that persist without ever reaching the singularity. Using Kruskal–Szekeres coordinates to avoid coordinate ambiguities, we identify these modes and analyze their eigenfunctions. Contrary to earlier claims, we find that the Regge-Wheeler master function of these modes have non-zero amplitude at the future event horizon but decay before interacting with the singularity. We consider observations of the modes along timelike geodesics. This work suggests that certain gravitational fluctuations can hover transiently within the black-hole interior, challenging common assumptions about wave behavior in uncharted and extreme regions of spacetime.
@@ I Introduction
Thus, studying the dynamical behavior of BHs is extremely important in order to achieve a robust understanding of how perturbations evolve in time and space. In BH exteriors, linear and non-linear excitations of asymptotically-flat and de Sitter geometries are bounded and decay with time (though grow with distance) [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], according to their spectral properties described by QNMs [79, 80, 81, 82, 83, 82, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 94, 97, 98, 99, 100, 101, 102, 103]. Quasibound states (QBSs) [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 115, 116, 117] of massive fields behave differently from QNMs due to different boundary conditions imposed at future null infinity. In fact, QBSs of massive fields, can grow or decay with time (and decay with distance). In BH interiors though, the dynamics of perturbations becomes more subtle and intricate. Actually, even though a potential spectrum in the interior of a BH might exist, we practically utilize it for purely theoretical reasons [118], in order to recognize the importance of singularities, realize the limitations of classical General Relativity (GR), and deduce models of quantum gravity [119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129] that can partially or completely erase singularities from BH interiors. All studies discussed above regarding the stability of BH interiors and Cauchy horizons in charged and rotating spacetimes share a common starting point; they begin with an initial value problem that is set in the exterior and evolves the initial data beyond the event horizon, and into the BH interior, where singularities lurk. A very important and systematic question is then posed: What is the behavior of perturbations that begin in the BH interior? Do they face their impending doom by plunging into the singularity or can they escape their inevitable fate?
@@
In this work, after analyzing the axial gravitational perturbations of the Schwarzschild BH interior, while assuming appropriate boundary conditions, we find overdamped quasibound states (OQBSs) which are localized inside the BH. This is achieved through the emergence of a negative well for axial gravitational perturbations in the BH interior, that diverges to $\infty$ close to the singularity. The respective wavefunctions of these OQBSs encounter the future event horizon but decay before encountering the singularity, or the past horizon, in the same manner as a QBS wavefunction would behave for $r\rightarrow\infty$ in the BH exterior, where $r$ is the usual radial coordinate in Schwarzschild coordinates. This analysis is achieved through the use of Kruskal–Szekeres coordinates [130, 131, 132] to avoid coordinate singularities and ambiguities, which has not yet been used for such a spectral problem. Our findings suggest the existence of special modes, the OQBSs, that hover indefinitely within the Schwarzschild BH’s interior.
@@
The spectrum of Schwarzschild BH interiors was also studied in Ref. [133] for scalar, axial and polar gravitational perturbations, but it was found that the internal modes did not encounter the event horizon, in contrast to our results. Ref. [133] also connected the internal modes with QNMs. Another study of the interior dynamics of a BH was made in Ref. [134], though the “spectral problem” solved in the interior used an inverted potential technique, as in Refs. [135, 136, 137], in order to find analytical bound states, through confluent Heun functions. Therefore, the bound state problem in [134] is fundamentally different from the one that we present here. In what follows we assume the geometrized unit system, i.e., $G=c=1$.
@@ II Schwarzschild black holes and perturbations
The Schwarzschild geometry is the unique asymptotically flat BH solution to the vacuum Einstein field equations in spherical symmetry [138]. The solution describes a BH with mass $M$, while its causal structure exhibits a null hypersurface at $r=2M$ that causally disconnects information from future null infinity when $r<2M$. On the other hand, when $r>2M$ information can reach future null infinity. Thus, the hypersurface at $r=2M$ separates the interior of the Schwarzschild BH from its exterior. Finally, at the BH interior there is a spacelike singularity at $r=0$, where the Kretschmann curvature scalar blows up. No observer can survive a spacelike singularity due to its infinite tidal forces and tidal deformations inflicted on it.
@@
The study of small perturbations in Schwarzschild BHs has a long history [139, 140]. Essentially, perturbations to the metric propagate as GWs on a fixed background; in our case the Schwarzschild background is described by the line element (1). Under the assumption of harmonic time dependence and by using the tensor spherical harmonics for the angular part’s eigenvalues, i.e. the spherical-harmonic index $\ell$, axial gravitational perturbations can be written in terms of a gauge-invariant master function $\psi(r)$ that satisfies the Regge-Wheeler equation [141, 142]
@@
Here, $\omega$ is the propagation frequency of the GW. The tortoise coordinate in Eq. (3) is defined as
@@
i.e. the sign changes inside the logarithm. In turn, in the interior we have $\infty<r_{*}<0$, where the singularity $r=0$ is mapped to $r_{*}=0$ and the event horizon $r=2M$ is mapped to $r_{*}\to$-$\infty$. Nevertheless, the Regge-Wheeler equation utilizes derivatives of the tortoise coordinate that still satisfy Eq. (5) inside the BH. In other words, the expression for $r_{*}$ is different inside and outside of the BH, but the expression for $dr_{*}$ is the same everywhere in the domain of $r_{*}$. Thus, Eqs. (3) and (6) are also valid in the Schwarzschild BH interior. We will take advantage of this, and attempt to set up a proper spectral problem in the interior of a Schwarzschild BH, that originates directly in the interior, i.e. initial data are given inside the BH for the evolution of small perturbations. We will choose the correct boundary conditions in the interior and find a spectrum of OQBSs.
@@ III Overdamped quasibound states inside a black hole
Fig. 1(a) demonstrates the effective potential Eq. (6) of axial gravitational perturbations both inside and outside the Schwarzschild BH. There is a weak maximum in the exterior of the BH ($r>2M$) which results in the well-known excitations of the light ring and the resulting QNMs [139]. An effective potential minimum appears in the BH interior, between the event horizon and the singularity, in both Fig. 1(a) which shows the potential with respect to $r$, as well as in Fig. 1(b) that shows the potential as a function of the tortoise coordinate $r_{*}$.
By Eq. (3), the internal minimum for axial gravitational perturbations can sustain bound states, each with negative effective energy $\omega^{2}$, as shown in the first row of Fig. 2. We focus on the simple case in which the effective energy is real. Thus, $\omega=i\omega_{\mathrm{I}}$ is purely imaginary, i.e., overdamped. Since only the square of $\omega$ appears in Eq. (3), the same spatial solution is obtained regardless of the sign of $\omega_{\mathrm{I}}$. This sign is determined by the ingoing boundary condition at the horizon. We consider the region of $r_{*}$ to the left of the potential minimum in Fig. 1(b), where $V_{\mathrm{RW}}$ is negligible in Eq. (3). The wavefunction $\psi$ should decay exponentially there for decreasing $r_{*}$, so $\psi\sim\mathrm{exp}(|\omega_{\mathrm{I}}|r_{*})$. Furthermore, the RW master function is given by $\Psi(t,r)=\mathrm{exp}(-i\omega t)\psi(r)=\mathrm{exp}(\omega_{\mathrm{I}}t)\psi(r)$, where $t$ is the Schwarzschild time coordinate. When $\omega_{\mathrm{I}}>0$, then $\Psi\sim\mathrm{exp}(\omega_{\mathrm{I}}v)$, where $v\equiv t+r_{*}$ the ingoing null coordinate. Since $\Psi$ is a function of $v$, it is ingoing at the horizon, which is physically required due to the fact that the event horizon is a one-way hypersurface where anything that reaches it inevitably fall into the BH. A similar calculation for $\omega_{\mathrm{I}}<0$ shows that $\Psi\sim\mathrm{exp}(\omega_{\mathrm{I}}u)$ where $u\equiv t-r_{*}$ is the outgoing null coordinate. There is no outgoing information from the BH interior whatsoever, so we must use $\omega_{\mathrm{I}}>0$ to satisfy the boundary condition at the event horizon.
On the other hand, since $r_{*}\sim-r^{2}/4M$ as $r\to 0$, the potential (6) diverges as $3/4r_{*}^{2}$ close to the singularity. Solving Eq. (3) analytically in the limit $r\rightarrow 0$ demonstrates that $\psi\sim(-r_{*})^{3/2}$. In summary, the internal boundary conditions for axial perturbations are
@@
We find the OQBSs by numerically solving Eq. (3) with the boundary conditions (8), in order to obtain the bound states of the axial effective potential in Eq. (6). We choose $\partial\psi\textfractionsolidus\partial r_{*}=1$ at some small starting value of $r_{*}$ in the range $10^{-5}<-r_{*}<3\times 10^{-4}$, and determine $\psi$ at the this point via the $(-r_{*})^{3/2}$ form of the boundary condition. Equation (3) is then integrated numerically from the starting point to some final negative value of $r_{*}$ which is chosen to be far enough from zero so that $\psi$ has decayed essentially to zero, as required from the boundary condition at the event horizon (8) for OQBSs. The frequency $\omega_{\mathrm{I}}$ is optimized to minimize $\psi^{2}$ at the final value of $r_{*}$ ($\psi$ is real). For the sake of quantifying the numerical noise in our scheme, the process is repeated 30 times for equally-spaced starting values of $r_{*}\in[10^{-5},3\times 10^{-4}]$.
@@ IV Results
The upper row of Figure 2 depicts the OQBS levels which are supported by the effective potential for each $\ell$. The values of $\omega_{\mathrm{I}}^{2}$ are indicated by colored horizontal lines. It is seen that larger choices of $\ell$ lead to deeper effective potential wells which support more energy states. The average and standard deviation of the resulting OQBSs of $\omega_{\mathrm{I}}$ are given in Table 1. The lower row of Figure 2 shows the eigenfunctions corresponding to each energy level.
@@
The number corresponding to the energy level of each OQBS is equal to the number of nodes in the wavefunction. The lowest, most negative (largest in absolute value) energy state has zero nodes, so it corresponds to the ground state $n=0$. It is obvious from Fig. 2 that excited states (shown with different colors) have more nodes, and thus define the excited OQBSs n=1, 2, 3n=1,\,2,\,3, and so on. The overall behavior of the eigenfunctions shown in Fig. 2 satisfy the boundary conditions (8), i.e., they tend to zero by a power law close to the singularity and decay exponentially at the event horizon, as a function of $r_{*}$. On the other hand, by using Kruskal-Szekeres coordinates, we will see that the full wavefunction encounters the horizon.
@@
To obtain a definitive picture on how an OQBS behaves at each point in spacetime, we transform the BH metric in study from Schwarzschild coordinates (t,r)(t,\,r) to Kruskal-Szekeres coordinates (U,V)(U,\,V) [130, 131, 132]. These coordinates are well-behaved everywhere outside the curvature singularity at $r_{*}=0$. Fig. 3 shows an OQBS in Kruskal-Szekeres coordinates, where deep blue indicates zero magnitude while brighter colors indicate larger values. We observe that the wavefunction vanishes at the singularity (dashed white curve at $UV=1$) and at the past event horizon (deep blue along $V=0$). We also find that the wavefunction fully encounters the future event horizon (brighter colors appear at $U=0$). Thus, the state is indeed quasi-bound, i.e., it is bound (zero) at the singularity but free (non-zero and finite) at the event horizon.
In order to determine the evolution of the mode according to a given observer, one should determine the path of the observer through the spacetime of Fig. 3. This approach avoids issues of Schwarzschild coordinates, such as the reversal of time and space inside the BH [134]. The red and green curves in Fig. 3 show examples of two such observers on timelike geodesics, each with almost constant $U$ or $V$. Thus, these two world lines almost form a light cone. Fig. 4 shows the values from Fig. 3 along the red and green world lines. The values are equal at the black circle, which corresponds to the white circle in Fig. 3, where the world lines happen to cross (see Appendix A for more details).
Figure 4: An OQBS as seen by geodesic observers. The red and green curves are the values from Fig. 3 along the red and green world lines, as functions of each observer’s proper time. The proper time is given in units of $M$. The two curves are equal at the black circle, which corresponds to the white circle in Fig. 3.
Finally, we point out that the isospectrality property between axial and polar metric perturbations is broken in the interior. In fact, it does not even hold for QBSs of massive fields at the exterior of Schwarzschild BHs [143, 144]. By breaking down the isospectrality problem into the exterior and interior of a Schwarzschild BH we conclude the following. In the exterior, axial and polar gravitational QNMs are isospectral. It has been shown analytically that the Regge-Wheeler (axial sector) and Zerilli (polar sector) equations are related via a transformation [139, 140], that was later associated with the Darboux transformation [149] and its generalized framework [150, 151, 152, 153], rendering the origin of the isospectrality more mathematically sound [154]. In addition to the perturbation invariance under the Darboux transformation, the boundary conditions make QNMs isospectral. In particular, the purely ingoing-wave condition at the event horizon and purely-outgoing wave condition at infinity are preserved under the Darboux transformation, therefore leading to isospectrality. The exterior spectral problem essentially shows that isospectrality is not only an outcome of the wave equation, but is also intrinsically and significantly tied to the boundary conditions. As we show in Appendix B, the regularity conditions at $r=0$ are not preserved under Darboux transformations and therefore the isospectrality of polar and axial metric perturbations is broken in the BH interior. This is similar to what happens in BHs in anti-de Sitter spacetime [155] for example, or even in exotic compact objects [156], that have different boundary conditions from the usual QNM ones at infinity and at the (would-be) horizon, and therefore do not satisfy any isospectrality relation.
Moving the discussion to the interior, and keeping in mind that the boundary conditions in Eq. (8) are not invariant under the Darboux transformation (at least the regularity condition close to the singularity is not satisfied, see Appendix B for a detailed proof), we can state that the interior axial OQBSs found here, will not satisfy any isospectrality relation with respect to the interior modes of polar perturbations, assuming that any polar QBSs might exist.

### v2

# Overdamped quasibound states inside a Schwarzschild black hole

@@ Abstract
Schwarzschild black-hole interiors, bounded by event horizons and terminated by spacelike singularities, are regions where all physical observers are inevitably destroyed. In the geometric optics approximation, waves follow null geodesics to the singularity. However, outside the geometric optics regime, the behavior of wave propagation can be rich and nuanced, even in such extreme habitats. In this work, we show that axial gravitational perturbations in the interior of a Schwarzschild black hole can form overdamped (non-oscillatory) quasibound states that decay before reaching the singularity. Using Kruskal-Szekeres coordinates to avoid coordinate ambiguities, we identify these modes and analyze their eigenfunctions. Contrary to earlier claims, we find that the Regge-Wheeler master function of these modes have non-zero amplitude at the future event horizon but decay before interacting with the singularity. We consider observations of the modes along timelike geodesics. This work suggests that certain gravitational fluctuations can hover transiently within the black-hole interior, challenging common assumptions about wave behavior in uncharted and extreme regions of spacetime.
@@ I Introduction
Thus, studying the dynamical behavior of BHs is extremely important in order to achieve a robust understanding of how perturbations evolve in time and space. In BH exteriors, linear and non-linear excitations of asymptotically-flat and de Sitter geometries are bounded and decay with time (though grow with distance) [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], according to their spectral properties described by QNMs [79, 80, 81, 82, 83, 82, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 94, 97, 98, 99, 100, 101, 102, 103]. Quasibound states (QBSs) [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 115, 116, 117] of massive fields behave differently from QNMs due to different boundary conditions imposed at future null infinity. In fact, QBSs of massive fields, can grow or decay with time (and decay with distance). In BH interiors though, the dynamics of perturbations becomes more subtle and intricate. Actually, even though a potential spectrum in the interior of a BH might exist, we practically utilize it for purely theoretical reasons [118], in order to recognize the importance of singularities, realize the limitations of classical General Relativity (GR), and deduce models of quantum gravity [119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129] that can partially or completely erase singularities from BH interiors. All studies discussed above regarding the stability of BH interiors and Cauchy horizons in charged and rotating spacetimes share a common starting point; they begin with an initial value problem that is set in the exterior and evolves the initial data beyond the event horizon, and into the BH interior, where singularities lurk. A very important and systematic question is then posed: What is the behavior of perturbations that begin in the BH interior? Do they simply plunge into the singularity, or can they exhibit a more complicated behavior?
@@
In this work, after analyzing the axial gravitational perturbations of the Schwarzschild BH interior, by imposing boundary conditions that are mathematically consistent with the geometry and the perturbation equations, we find purely-damped (non-oscillatory) modes that are localized solely in the internal region of the BH. We name these solutions as overdamped quasibound states (OQBSs). These, otherwise, regular solutions are sourced through the emergence of a negative well for axial gravitational perturbations in the BH interior, that diverges to $\infty$ close to the singularity. The respective wavefunctions of these OQBSs encounter the future event horizon but decay before encountering the singularity, or the past horizon, in the same manner as a QBS wavefunction would behave for $r\rightarrow\infty$ in the BH exterior, where $r$ is the usual radial coordinate in Schwarzschild coordinates. In order to avoid certain ambiguities regarding the nature of Schwarzschild coordinates inside the BH, and bypass the coordinate singularity at the event horizon, we have achieved this analysis through the use of Kruskal–Szekeres coordinates [130, 131, 132], which have not yet been used for such a spectral problem. Our findings suggest that the OQBSs are regular everywhere in the Schwarzschild interior, hover transiently and eventually decay with respect to proper time, before encountering the curvature singularity.
@@
The spectrum of Schwarzschild BH interiors was also studied in Ref. [133] for scalar, axial and polar gravitational perturbations, but it was found that the internal modes did not encounter the event horizon, in contrast to our results. Ref. [133] also connected the internal modes with asymptotic QNMs. Another study of the interior dynamics of a BH was performed in Ref. [134], though the “spectral problem” solved in the interior used an inverted potential technique, as in Refs. [135, 136, 137], in order to find analytical bound states, through confluent Heun functions. Therefore, the bound state problem in [134] is fundamentally different from the one that we present here. In what follows we assume the geometrized unit system, i.e., $G=c=1$.
@@ II Schwarzschild black holes and perturbations
The Schwarzschild geometry is the unique asymptotically flat BH solution to the vacuum Einstein field equations in spherical symmetry [138]. The solution describes a BH with mass $M$, while its causal structure exhibits a null hypersurface at $r=2M$ that causally disconnects information from future null infinity when $r<2M$. On the other hand, when $r>2M$ information can reach future null infinity. Thus, the hypersurface at $r=2M$ separates the interior of the Schwarzschild BH from its exterior. Finally, at the BH interior there is a spacelike curvature singularity at $r=0$, where the Kretschmann scalar blows up. No observer can survive a spacelike singularity due to its infinite tidal forces and tidal deformations inflicted on it.
@@
The study of small perturbations in Schwarzschild BHs has a long history [139, 140]. Essentially, perturbations to the metric propagate as GWs on a fixed background; in our case the Schwarzschild background is described by the line element (1). Under the assumption of harmonic dependence in the Schwarzschild coordinate $t$ and by using the tensor spherical harmonics for the angular part’s eigenvalues, i.e. the spherical-harmonic index $\ell$, axial gravitational perturbations can be written in terms of a gauge-invariant master function $\psi(r)$ that satisfies the Regge-Wheeler equation [141, 142]
@@
Here, $\omega$ is the propagation frequency of the axial GW. The tortoise coordinate in Eq. (3) is defined as
@@
i.e. the sign changes inside the logarithm. In turn, in the interior we have $\infty<r_{*}<0$, where the singularity $r=0$ is mapped to $r_{*}=0$ and the event horizon $r=2M$ is mapped to $r_{*}\to$-$\infty$. Nevertheless, the Regge-Wheeler equation utilizes derivatives of the tortoise coordinate that still satisfy Eq. (5) inside the BH. In other words, the expression for $r_{*}$ is different inside and outside of the BH, but the expression for $dr_{*}$ is the same everywhere in the domain of $r_{*}$. Thus, Eqs. (3) and (6) are also valid in the Schwarzschild BH interior. We will take advantage of this, and attempt to set up a proper spectral problem in the interior of a Schwarzschild BH, that originates directly in the interior, i.e. initial data are given inside the BH for the evolution of small perturbations. Our analysis is directly based on the Regge–Wheeler equation, which describes axial gravitational perturbations in the Schwarzschild geometry. While inside the event horizon $r$ and $r_{*}$ are timelike, this does not affect the validity of the equation itself, which remains mathematically well-defined in both the exterior and interior of the BH spacetime. In what follows, we retain the conventional terminology of the “tortoise coordinate”, in order to remain consistent with the literature. By choosing appropriate boundary conditions in the interior we find the spectrum of OQBSs. This approach was used in previous work to compute QNMs in the interior of other BH geometries [118, 149]. However, even though the initial calculations are performed in Schwarzschild coordinates, we ultimately use the Kruskal-Szekeres coordinate system in order to interpret the physical character of these modes.
@@ III Overdamped quasibound states inside a black hole
Fig. 1(a) demonstrates the effective potential Eq. (6) of axial gravitational perturbations both inside and outside the Schwarzschild BH. There is a weak maximum in the exterior of the BH ($r>2M$) which results in the well-known excitations of the light ring and the resulting QNMs [139]. An effective potential minimum appears in the BH interior, between the event horizon and the singularity, in both Fig. 1(a) which shows the potential with respect to $r$, as well as in Fig. 1(b) that shows the potential as a function of the “tortoise coordinate” $r_{*}$.
The interior of a BH is often described in the geometric optics limit, in which GWs behave as massless particles which follow null geodesics [150, 151]. The waves are described by Eq. 3, in which $\omega^{2}$ plays the role of an effective energy. If $\omega$ is sufficiently large, $V_{\mathrm{RW}}$ becomes negligible and one obtains the geometric optics limit, as illustrated in Fig. 1(b). Restoring the time derivative, Eq. 3 becomes
$\frac{\partial^{2}\psi}{\partial r_{*}^{2}}-\frac{\partial^{2}\psi}{\partial t^{2}}\sim 0,$ (8)
for asymptotically large $\omega$. This equation is solved by any function of $t\pm r_{*}$. In particular, it is solved by $\delta\left(t+r_{*}\right)$ and $\delta\left(t-r_{*}\right)$, which correspond to pointlike massless particles following null geodesics.
In this work, we are interested in wave behavior which is not within the geometric optics limit. By Eq. (3), the internal minimum for axial gravitational perturbations can sustain bound states, each with negative effective energy $\omega^{2}$, as shown in the first row of Fig. 2. We focus on the simple case in which the effective energy is real. Thus, $\omega=i\omega_{\mathrm{I}}$ is purely imaginary, i.e., overdamped. Since only the square of $\omega$ appears in Eq. (3), the same spatial solution is obtained regardless of the sign of $\omega_{\mathrm{I}}$. This sign is determined by the ingoing boundary condition at the horizon. We consider the region of $r_{*}$ to the left of the potential minimum in Fig. 1(b), where $V_{\mathrm{RW}}$ is negligible in Eq. (3). The wavefunction $\psi$ should decay exponentially there for decreasing $r_{*}$, so $\psi\sim\mathrm{exp}(|\omega_{\mathrm{I}}|r_{*})$. Furthermore, the RW master function is given by $\Psi(t,r)=\mathrm{exp}(-i\omega t)\psi(r)=\mathrm{exp}(\omega_{\mathrm{I}}t)\psi(r)$, where $t$ is the Schwarzschild time coordinate. When $\omega_{\mathrm{I}}>0$, then $\Psi\sim\mathrm{exp}(\omega_{\mathrm{I}}v)$, where $v\equiv t+r_{*}$ the ingoing null coordinate. Since $\Psi$ is a function of $v$, it is ingoing at the horizon, which is physically required due to the fact that the event horizon is a one-way hypersurface where anything that reaches it inevitably fall into the BH. A similar calculation for $\omega_{\mathrm{I}}<0$ shows that $\Psi\sim\mathrm{exp}(\omega_{\mathrm{I}}u)$ where $u\equiv t-r_{*}$ is the outgoing null coordinate. There is no outgoing information from the BH interior whatsoever, so we must use $\omega_{\mathrm{I}}>0$ to satisfy the boundary condition at the event horizon.
On the other hand, since $r_{*}\sim-r^{2}/4M$ as $r\to 0$, the potential (6) diverges as $3/4r_{*}^{2}$ close to the singularity. In this limit, Eq. (3) has a bound analytic solution of the form $\psi\sim(-r_{*})^{3/2}$. In summary, the internal boundary conditions for axial perturbations are
@@
We find the OQBSs by numerically solving Eq. (3) with the boundary conditions (9), in order to obtain the bound states of the axial effective potential in Eq. (6). We choose $\partial\psi\textfractionsolidus\partial r_{*}=1$ at some small starting value of $r_{*}$ in the range $10^{-5}<-r_{*}<3\times 10^{-4}$, and determine $\psi$ at this point via the $(-r_{*})^{3/2}$ form of the boundary condition. Equation (3) is then integrated numerically from the starting point to some final negative value of $r_{*}$ which is chosen to be far enough from zero so that $\psi$ has decayed essentially to zero, as required from the boundary condition at the event horizon (9) for OQBSs. The frequency $\omega_{\mathrm{I}}$ is optimized to minimize $\psi^{2}$ at the final value of $r_{*}$ ($\psi$ is real). For the sake of quantifying the numerical noise in our scheme, the process is repeated 30 times for equally-spaced starting values of $r_{*}\in[10^{-5},3\times 10^{-4}]$.
@@ IV Results
The upper row of Fig. 2 depicts the OQBS levels which are supported by the effective potential for each $\ell$. The values of $\omega_{\mathrm{I}}^{2}$ are indicated by colored horizontal lines. It is seen that larger choices of $\ell$ lead to deeper effective potential wells which support more energy states. The average and standard deviation of the resulting OQBSs of $\omega_{\mathrm{I}}$ are given in Table 1. The lower row of Fig. 2 shows the eigenfunctions corresponding to each energy level. The number corresponding to the energy level of each OQBS is equal to the number of nodes in the wavefunction. The lowest, most negative (largest in absolute value) energy state has zero nodes, so it corresponds to the ground state $n=0$. It is obvious from Fig. 2 that excited states (shown with different colors) have more nodes, and thus define the excited OQBSs n=1, 2, 3n=1,\,2,\,3, and so on. The overall behavior of the eigenfunctions shown in Fig. 2 satisfy the boundary conditions (9), i.e., they tend to zero by a power law close to the singularity and decay exponentially at the event horizon, as a function of $r_{*}$. Nevertheless, in order to fully grasp the physical interpretation of these modes, we switch to Kruskal-Szekeres coordinates, which unveils that the full wavefunction encounters the event horizon.
@@
To obtain a definitive picture on how an OQBS behaves at each point in spacetime, we transform the BH metric in study from Schwarzschild coordinates (t,r)(t,\,r) to Kruskal-Szekeres coordinates (U,V)(U,\,V) [130, 131, 132]. These coordinates are well-behaved everywhere outside the curvature singularity at $r_{*}=0$. Fig. 3 shows an OQBS in Kruskal-Szekeres coordinates, where deep blue indicates zero magnitude while brighter colors indicate larger values. We observe that the wavefunction vanishes at the singularity (dashed white curve at $UV=1$) and at the past event horizon (deep blue along $V=0$). We also find that the wavefunction fully encounters the future event horizon (brighter colors appear at $U=0$). Thus, the state is formally quasibound, i.e., it decays (to zero) at the singularity but it is free (non-zero and finite) at the event horizon.
In order to determine the evolution of the mode according to a given observer, one should determine the path of the observer through the spacetime of Fig. 3. This approach avoids issues of Schwarzschild coordinates, such as the reversal of time and space inside the BH [134]. The red and green curves in Fig. 3 show examples of two such observers on timelike geodesics, each with almost constant $U$ or $V$. Thus, these two world lines almost form a light cone. Fig. 4 shows the values from Fig. 3 along the red and green world lines. The values are equal at the black circle, which corresponds to the white circle in Fig. 3, where the world lines happen to cross (see Appendix A for more details). Figure 4 ultimately demonstrates that the OQBS decays with proper time for both observers as they approach the singularity, in accord with the boundary conditions. Furthermore, oscillations are seen in Fig. 4 as a function of proper time, which reflect the oscillations as a function of $r_{*}$ seen in Fig. 2, while there are no oscillations as a function of the coordinate time $t$. This association of proper time with $r_{*}$ is expected inside a black hole, where the roles of space and time are interchanged.
It is interesting to ask whether the OQBSs produce finite metric perturbations at the singularity. In the geometric optics limit illustrated in Fig. 1(b) [150, 151], $V_{\mathrm{RW}}$ is negligible compared to $\omega^{2}$ in Eq. (3) (except near the singularity), and the gravitational perturbations can come very close to the singularity, although they would not reach it due to the infinite potential barrier at $r=0$. The OQBS wavefunction vanishes at the singularity like $\psi\sim(-r_{*})^{3/2}$, and the radial part of the axial metric perturbation is related to the quantity $r\,\psi(r)\sim r_{*}^{2}$ [140], so it tends to zero at $r=0$ (see also Eq. (B.21) in [152]). Therefore, the OQBSs produce vanishing metric perturbations at the singularity $r=0$.
Finally, we point out that the isospectrality property between axial and polar metric perturbations seems to be broken in the interior. It certainly does not hold for QBSs of massive fields at the exterior of Schwarzschild BHs [143, 144]. By breaking down the isospectrality problem into the exterior and interior of a Schwarzschild BH we conclude the following. In the exterior, axial and polar gravitational QNMs are isospectral. It has been shown analytically that the Regge-Wheeler (axial sector) and Zerilli (polar sector) equations are related via a transformation [139, 140], that was later associated with the Darboux transformation [153] and its generalized framework [154, 155, 156, 157], rendering the origin of the isospectrality more mathematically sound [158]. In addition to the perturbation invariance under the Darboux transformation, the boundary conditions make QNMs isospectral. In particular, the purely ingoing-wave condition at the event horizon and purely-outgoing wave condition at infinity are preserved under the Darboux transformation, therefore leading to isospectrality. The exterior spectral problem essentially shows that isospectrality is not only an outcome of the wave equation, but is also intrinsically and significantly tied to the boundary conditions. As we show in Appendix B, the regularity conditions at $r=0$ are not preserved under Darboux transformations and therefore the isospectrality of polar and axial metric perturbations might be broken in the BH interior. This is similar to what happens in BHs in anti-de Sitter spacetime [159] for example, or even in exotic compact objects [160], that have different boundary conditions from the usual QNM ones at infinity and at the (would-be) horizon, and therefore do not satisfy any isospectrality relation.
Moving the discussion to the interior, and keeping in mind that the boundary conditions in Eq. (9) are not invariant under the Darboux transformation (at least the regularity condition close to the singularity is not satisfied, see Appendix B for a detailed proof), we can assume that the interior axial OQBSs found here, may not satisfy the isospectrality relation with respect to the interior modes of polar perturbations.